OpenStudy (anonymous):

chemical name of Pb(BrO2)2?

OpenStudy (anonymous):

BrO2 is an oxyanion, and the rules for those are that the most common one takes the name of the element other than oxygen and adds "-ate," so the most common oxyanion of bromine (which is BrO3-) is named "bromate." If you subtract one oxygen, you change "ate" to "ite", so this anion is named "bromite." (If you subtract another oxygen, you get BrO-, hypobromite, and if you add one you get BrO4-, perbromate.) The BrO2- anion has a charge of -1. (How do I know that? All the oxyanions using elements of Group 7A have charge -1. Those made from elements in Group 6A have charge -2, and those using elements from Group 5A have charge -3. See a pattern? Unfortunately it breaks in Group 4A and 3A, because carbonate CO3-2 has charge -2 instead of -4, and borate BO3-3 has charge -3 instead of -5.) With two bromites, you need a +2 to balance and make each formula unit neutral. That means the lead cation must be Pb+2. It's name will be "lead(II)," with the Roman numeral II indicating the +2 charge. So the full name is "lead(II) bromite." Similar examples: Pb(BrO)2 = lead(II) hypobromite Pb(BrO2)2 = lead(IV) bromite PbCO3 = lead(II) carbonate

OpenStudy (anonymous):

Lead(II) Bromite

OpenStudy (anonymous):

Bromite